OpenStudy (anonymous):

Answer for fan + best answer. Estimate by rounding to the nearest whole number. twenty-four and two-elevenths divided by three and six-sevenths 4 6 8 twelve

OpenStudy (goldphenoix):

\[24 \frac{ 2 }{ 11 } \div 3 \frac{ 6 }{ 11 }\] You can solve this by doing the method KCF. (Keep, Change, Flip.) But before we do that, we must change the mixed number into improper fraction. You can do this by multiplying the denominator from from the whole and then adding it from the numerator, but the denonimator stay the same. So it should look: \[\frac{ 266 }{ 11 } \div \frac{ 39 }{ 11 }\] Now we can use the method, KCF!!! So you keep the first fraction. We change the division sign into a multiplication sign. Then we flip the second fraction. Now it should look like: \[\frac{ 266 }{ 11 } \times \frac{ 11 }{ 39 }\] The final step is to multiply across. 266 x 11 = 2926. 11 x 39 = 429. \[\frac{ 2926 }{ 429 }\] Okay there's one more step. You want to simplify it. You can simplify it by dividing the numerator from the denonimator. This should give you 6.82. Round up to 7.

OpenStudy (goldphenoix):

Apparently there's no 7, let me check my work to see if I did anything wrong.

OpenStudy (anonymous):

Yeah there isn't, thanks.

OpenStudy (goldphenoix):

I checked it. I even used a calculator. It's ~6.82.

OpenStudy (goldphenoix):

It's either 6 or 8, but I recommend choosing 6. :)

OpenStudy (anonymous):

hmm. I'll just go with 6. Thanks!

OpenStudy (goldphenoix):

I hope this help! Have a nice day/afternoon/night. :)